OpenStudy (anonymous):

Rationalize the denominator \[\frac{sqrt{2a}}{sqrt{a^2+1}}\]

OpenStudy (anonymous):

\[=\frac{\sqrt{2a}}{\sqrt{a^2+1}}\] \[=\frac{\sqrt{2a}}{\sqrt{a^2+1}}*\frac{\sqrt{a^2+1}}{\sqrt{a^2+1}}\]

OpenStudy (anonymous):

\[\frac{\sqrt{2a(a^2+1)}}{\sqrt{a^2+1}}\] \[\frac{\sqrt{2a^3+2a}}{a^2+1}\]

OpenStudy (anonymous):

I don't understand the conjugate

OpenStudy (mertsj):

What would be the conjugate of sqrt(a^2+1)?

OpenStudy (anonymous):

I don't know

OpenStudy (mertsj):

The conjugate is a binomial that results in an (a+b)(a-b) situation.

OpenStudy (mertsj):

Exactly. There isn't one.

OpenStudy (anonymous):

It's a sum of square we didn't get into that territory

OpenStudy (anonymous):

So you multiply it by itself?

OpenStudy (mertsj):

Just like you did because that rationalizes it. And that was the requirement of the problem. The reason you multiply by the conjugate (in appropriate circumstances) is because it rationalizes some desired quantity.

OpenStudy (anonymous):

For this particular problem, you multiply by itself because that would get rid of the radical, but usually, you multiply by the conjugate (difference in squares) to get rid of the middle term which would contain the radical.

OpenStudy (anonymous):

So if it was \[\sqrt{x^2-1}\] I would multiply by conjugate?

OpenStudy (mertsj):

NO!! That does not have a conjugate.

OpenStudy (mertsj):

\[\frac{2}{3-\sqrt{5}}\]

OpenStudy (anonymous):

Only if it's seperate terms it has a conjugate? Okay. Thanks so much everyone!!!

OpenStudy (anonymous):

Here's an example of when you WOULD use a conjugate: \[\frac{\sqrt{2a}}{\sqrt{a} + \sqrt{b}}\]

OpenStudy (mertsj):

There's a case where you would multiply by the conjugate. Do you see that it is a binomial? And the conjugate is a binomial?

OpenStudy (mertsj):

And, my dear, please learn that there is A RAT in sepARATe

OpenStudy (anonymous):

lol. SepErate: Verb SepArate: Adjective

OpenStudy (anonymous):

Oh, lol... sorry @mertsj thanks once again!